The US athlete Florence Griffith-Joyner won the 100 m sprint gold medal at Seol Olympic 1988 setting a new Olympic record of 10.54 s. Assume that she achieved her maximum speed in a very short-time and then ran the race with that speed till she crossed the line. Take her mass to be 50 kg. (a) Calculate the kinetic energy of Griffith-Joyner at her full speed. (b) Assuming that the track, the wind etc. offered an average resistance of one tenth of her weight, calculate the work done by the resistance during the run. (c) What power Griffith-Joyner had to exert to maintain uniform speed?

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. S = 100 m, t = 10.54 sec, m = 50 kg The motion can be assumed to be uniform because the time taken for acceleration is minimum. a) Speed v = S/t = 9.487 e/s So, K.E. = ½ mv^2 = 2250 J b) Weight = mg = 490 J given R = mg /10 = 49 J so, work done against resistance WF = – RS = – 49 × 100 = – 4900 J c) To maintain her uniform speed, she has to exert 4900 j of energy to over come friction P = W/t = 4900 / 10.54 = 465 W